A lump sum of money is invested at a rate of 10% per annum compounded quarterly. How 

long will it take for the investment to triple?


1
Expert's answer
2022-02-11T11:11:55-0500

"P=A(1+\frac{0.10}{4})^{4t}" .

If P=3A: "3=(1.025)^{4t}" .

"4tln{1.025}=ln3" .

"t=\frac{ln3}{4ln1.025}\approx 11.12\;yr=133.44\;months=4059\;days."